Triangle ABC and Triangle RST are similar with m < A < R and m < B = m < S. If AB, BC, and AC are 6 inches, 8 inches and 9 inches respectively, and RS is 7.2 inches, find ST (Answer to the nearest tenth)
Answer by midwood_trail(310)

AB/RS = BC/ST
We are looking for ST.
6/7.2 = 8/ST
Cross-multiply and then divide.
7.2 X 8 = 6 X ST
57.6 = 6ST
57.6/6 = ST
9.6 = ST